### v5

# Ivy-Fake: A Unified Explainable Framework and Benchmark for Image and Video AIGC Detection

Abstract
The rapid development of Artificial Intelligence Generated Content (AIGC) techniques has enabled the creation of high-quality synthetic content, but it also raises significant security concerns. Current detection methods face two major limitations: (1) the lack of multidimensional explainable datasets for generated images and videos. Existing open-source datasets (e.g., WildFake, GenVideo) rely on oversimplified binary annotations, which restrict the explainability and trustworthiness of trained detectors. (2) Prior MLLM-based forgery detectors (e.g., FakeVLM) exhibit insufficiently fine-grained interpretability in their step-by-step reasoning, which hinders reliable localization and explanation. To address these challenges, we introduce Ivy-Fake, the first large-scale multimodal benchmark for fake image and video detection. It consists of over 106K richly annotated training samples (images and videos) and 5,000 manually verified evaluation examples, sourced from multiple generative models and real-world datasets through a carefully designed pipeline to ensure both diversity and quality. Furthermore, we propose Ivy-xDetector, a multimodel large language model (MLLM) based on reinforcement fine-tuning (RFT), capable of producing explainable reasoning chains and achieving robust performance across multiple fake image and video detection benchmarks.
[…]
1 Introduction
Figure 1: Overview of the Ivy-Fake framework: By conducting in-depth analysis of temporal and spatial artifacts, the framework enables explainable detection of AI-generated content.
The rapid development of diffusion-based models has triggered an exponential growth in Artificial Intelligence Generated Content (AIGC), like Sora (sora_videoworldsimulators2024), DALL-E (dalle_3_betker2023improving), Imagen (imagegen_saharia2022photorealistic), and Stable Diffusion (rombach2022high), which have redefined state-of-the-art performance in text-to-image synthesis. However, these advances also raise significant security concerns, including the misuse of DeepFake (Qureshi2024DeepfakeFA), document tampering (Qureshi2024DeepfakeFA; qu2023doctamper; qu2024miml; qu2025ostf), and dataset poisoning (10.5555/3294996.3295110; ying2026beyond). The increasing realism of synthetic content blurs the boundary between genuine and fabricated media, posing critical challenges for misinformation control, content provenance verification, and the preservation of public trust.
However, most existing approaches primarily focus on binary authenticity classification (demamba; wildfake_2025; aigvdet_bai2024ai; aide_yan2025a), which limits the human interpretability of model predictions (qu2026textshield; qu2024omni; loki_ye2024loki; fakehr1). Existing benchmarks also remain inadequate for evaluating explainable AIGC detection. Datasets such as AIGCDetectionBenchmark (aigcdetectionbenchmark_zhong2023patchcraft) and GenVideo (demamba) provide only binary labels, while more recent resources like LOKI (loki_ye2024loki) attempt to incorporate fine-grained anomaly annotations across modalities but are still constrained in scale and diversity. On the other hand, current methods such as AIDE (aide_yan2025a) and Demamba (demamba) are limited to a single modality—either image or video—without exploring unified detection across both domains. Similarly, existing datasets exhibit the same fragmentation: FakeBench (liu2024mmfakebench) emphasizes explainable fake image detection but omits video content, whereas FakeClue (fakeclue_wen2025spot) provides extensive image-level annotations yet lacks integrated video data. This fragmentation leads to substantial gaps in benchmarking and hinders unified advancement in explainable multimodal AIGC detection.
To address these challenges, we propose Ivy-Fake, a comprehensive dataset designed to evaluate explainable multimodal AIGC detection. Ivy-Fake offers: 1) Diverse Multimodal Data, a large-scale dataset comprising 61,107 annotated images and 45,272 annotated videos for training, along with 5,000 samples for evaluation. 2) Explainable Annotations, rich annotations that extend beyond binary labels to include detailed reasoning, enabling nuanced evaluation of models’ interpretability and explanatory capabilities.
In addition to providing Ivy-Fake, we introduce Ivy-xDetector, a model for detecting AI-generated images and videos and explaining the associated artifacts. Unlike other MLLMs, Ivy-xDetector excels at spotting generative artifacts, spatial in images and both spatial and temporal in videos. By integrating multiple spatial and temporal feature extractors, it detects image-level artifacts and video-level temporal inconsistencies with superior accuracy. Through the incorporation of explainable annotations and carefully curated domain-diverse data, our approach achieves state-of-the-art performance using only a fraction of the data required by existing detectors. While prior image and video classifiers (demamba; aide_yan2025a; aigcdetectionbenchmark_zhong2023patchcraft) often rely on millions (M-level) of samples, our method attains superior performance across both image and video AIGC detection tasks with fewer than 200K samples.
Our main contributions are summarized as follows:
  • We introduce Ivy-Fake, the first large-scale dataset for explainable AIGC detection across both images and videos, comprising 106,379 training samples and 5,000 manually verified test instances. Each entry is enriched with fine-grained visual annotations and textual reasoning to support transparent multimodal evaluation and future research.
  • We propose Ivy-xDetector, a reinforcement learning–based model by Reinforcement Learning with Verifiable Rewards (RLVR) built upon Group Relative Policy Optimization (GRPO), which achieves superior performance on multiple image and video detection benchmarks using fewer than 200K samples, significantly outperforming existing methods that rely on millions of data points.
  • We conduct extensive experiments to validate the effectiveness of both the dataset and the model, demonstrating their substantial impact on improving multimodal large models’ capability in synthetic content detection.
Table 1: Task examples of Ivy-Fake.
Features Sub Dimension Example
Spatial Impractical Which visual clue reveals an impossible spatial or lighting configuration in the scene?
Luminosity (A) Light direction opposite source. (B) Face without lighting. (C) Dark background, bright face. (D) Bright background, dark face.
Localized Which clue shows local blur or focus inconsistency?
Blur (A) Clear face but blurred hand. (B) Sharp background, fuzzy hair. (C) Glasses edges blurred, eyes clear. (D) Hand blurred while holding clear object.
Illegible Which part of the document text is tampered or altered?
Letters (A) Text blurred or half missing. (B) Letters inconsistent in size. (C) Words overlap or melt together. (D) Signs contain random symbols.
Omitted Which part of the person looks incomplete or missing?
Components (A) Missing one finger. (B) Two fingers merged. (C) Missing eyebrows. (D) Three pairs of glasses.
Spatial Where does the spatial relation look unnatural or physically impossible?
Relationships (A) Arm passes through body. (B) Legs overlap unnaturally. (C) Head detached from neck. (D) Hand holds object from wrong side.
Distorted Which part appears deformed or stretched in shape?
Components (A) Finger bent at odd angle. (B) Eye stretched sideways. (C) Mouth warped or uneven. (D) Hand larger than face.
Chromatic Which clue shows unnatural or inconsistent colors in the image?
Irregularity (A) Face partly blue under warm light. (B) Green reflection on red shirt. (C) Sky color bleeds into hair. (D) Background color unevenly saturated.
Abnormal What visual area shows abnormal or inconsistent texture?
Texture (A) Skin looks like plastic. (B) Hair pattern repeats unnaturally. (C) Cloth surface too smooth or glossy. (D) Wall texture mixed with skin.
Temporal Luminance Which moment shows sudden or inconsistent brightness between frames?
Discrepancy (A) Light flickers without reason. (B) Face turns dark while background stays bright. (C) Shadow jumps between frames. (D) Object brightness changes abruptly.
Awkward When does the facial movement look unnatural or discontinuous across frames?
Facial Expression (A) Smile appears and disappears suddenly. (B) Eyes blink in different directions. (C) Mouth freezes mid-motion. (D) Expression changes without emotion cue.
Duplicated Where does an object or body part appear twice across frames?
Component (A) Two faces shown in one frame. (B) Hand shadow moves separately. (C) Extra arm appears then fades. (D) Same object duplicated in next frame.
Non-Spatial When does the action sequence look illogical or out of order?
Relationships (A) Person closes door before touching it. (B) Cup moves without contact. (C) Smile appears after laughter. (D) Object vanishes then reappears suddenly.
Figure 2: Token Length Distributions and Multi-Dimensional Coverage Across Datasets; Left: Distribution of token lengths across datasets; Right: Coverage of multiple dimensions in explainability datasets, extracted using Qwen3-32B (qwen3). The Prompt can be seen in appendix.
2 Related Work
2.1 Methods for Synthetic Content Detection
Due to growing concerns about the misuse of synthetic data (deng2024survey; qu2025ostf; li2025artificial), research on AI-generated content (AIGC) detection has expanded rapidly in recent years (wang2026forgeryvcr; park2025vidguard; shuai2026detectors; tan2026videoveritas; tan2025veritas; yu2026agentfox; huang2025unishield; Zhou2025AIGIHolmesTE). Most existing models for AI-generated images and videos formulate the task as binary classification, simply predicting whether the content is "real" or "fake." Representative examples include CNN-based AIGVDet (aigvdet_bai2024ai), CNNSpot (cnnspot_wang2020cnn), DIRE (dire_wang2023dire), Mirror (liu2026mirror) and AIDE (aide_yan2025a). Meanwhile, several works have explored the application of multimodal large language models (MLLMs) to AIGC detection, including Synartifact (cao2024synartifact) and Bi-LORA (keita2025bi). However, these approaches largely overlook the importance of interpretability in AIGC detection.
Some efforts attempt to introduce interpretability by leveraging spatial annotations (qu2023doctamper; zhu2025mesorch; qu2024miml; qu2024omni), frequency-domain artifact analysis (zhang2023perceptual) or or hybrid attention (xin2026hytrechybridtemporalawareattention; kong2025token). Nevertheless, the resulting explanations are often difficult for humans to comprehend, as they lack clarity in natural language. This limitation is particularly evident in the video domain, where AI-generated content frequently exhibits obvious flaws, e.g., incoherent frame transitions and object inconsistency, that are easily noticed and reasoned about by humans (deng2024survey). FakeClue (fakeclue_wen2025spot) introduces the use of vision-language models (VLMs) to provide interpretability for image-level detection, but it does not offer a unified framework that integrates both images and videos.
2.2 Datasets for Synthetic Content Detection.
Early datasets for synthetic content detection (zhang2025detecting; jiang2025revisiting; ma2025imdl; du2025forensichub; zhu2025does), such as CNNSpot (cnnspot_wang2020cnn), primarily collected fake images generated by GAN-based models (gan1_goodfellow2014generative; gan2_zhu2017unpaired; gan3_brock2018large). However, with the advent of more advanced generative architectures like diffusion models (diff4_hertz2022prompt; diff5_nichol2021glide; ying2026beyond) and their variants, the authenticity of generated content has significantly increased, making it more challenging for detection models to discern. This has spurred the development of newer datasets, including ArtiFact (cao2024synartifact), GenImage (genimage_2023), GenImagePP (zhou2025breaking) and WildFake (wildfake_2025). GenImage (genimage_2023) comprises images from the 1000 ImageNet (imagenet_ILSVRC15) categories, generated by eight state-of-the-art generators such as Stable Diffusion (rombach2022high) and Midjourney. Nevertheless, these datasets predominantly focus on image-based content. More recently, datasets emphasizing interpretability have also been introduced (zhao2026spinebench). FakeClue (fakeclue_wen2025spot) contains a large amount of image data with explainability annotations but lacks video data. LOKI (loki_ye2024loki) offers data across 26 different categories and includes 18,000 distinct questions; however, its volume of multimodal data is relatively small and primarily suited for evaluation rather than comprehensive model training (ji2025zoom). Therefore, a critical gap exists for a unified benchmark encompassing both image and video modalities to rigorously evaluate the performance of contemporary AIGC detectors.
3 Methodology
3.1 Data Collection
The Ivy-Fake dataset is curated from GenVideo (demamba), LOKI (loki_ye2024loki), FakeClue (fakeclue_wen2025spot), WildFake (wildfake_2025), Kinetics-400 (kay2017kineticshumanactionvideo), and MMFakeBench (liu2024mmfakebench). Since Ivy-Fake is substantially larger than other datasets, we primarily detail our data processing procedures here.
Image/Video Corruption Filtering.
The integrity of images and videos is critical for MLLM training. We observed that the official GenVideo (demamba) repository contained a large number of unreadable videos. To ensure data quality, we employed the official Qwen2.5-VL (bai2025qwen2) I/O library to filter out corrupted files, discarding images and videos that failed to load.
Image/Video Resolution.
To maintain consistency, we only retained images whose resolutions fell within [3,136, 12,845,056] pixels, and videos whose per-frame resolutions were within [100,352, 602,112] pixels, with the additional constraint that total pixels did not exceed 19,267,968. Samples outside these ranges were discarded.
Near-Duplicate Image Filtering.
Selecting challenging and diverse samples is crucial for model robustness. We extracted visual features from images and video frames using CLIP (Radford2021LearningTV) and computed intra-class similarity scores. Images with excessively high similarity were removed, while videos were left unprocessed. To comprehensively activate and enhance the model’s reasoning capabilities across the task, we propose a high-quality data construction pipeline that integrates explainable detection of AI-generated content.
Public Dataset Sources
The Ivy-Fake training set comprises approximately 106K image and video samples collected during the cold-start phase from diverse sources to ensure broad coverage of contemporary generative models. A significant portion was sourced from established public datasets, including GenVideo (demamba), LOKI (loki_ye2024loki), FakeClue (fakeclue_wen2025spot), WildFake (wildfake_2025), Kinetics-400 (kay2017kineticshumanactionvideo), MMFakeBench (liu2024mmfakebench), GenImage (genimage_2023), DiffusionForensics (dire_wang2023dire). The cold-start phase aimed to enhance the granularity of interpretable content. To further improve the detector’s accuracy, an additional 87,500 image/video samples were stratified and sampled from multiple datasets for RLVR training, including MSRVTT (xu2016msrvtt) (9,000 videos), GenImage (genimage_2023) (32,000 images per category), CnnSpot (cnnspot_wang2020cnn) (19,000 samples), DigiFakeAV (liu2025faceswappingdiffusionbaseddigital) (2,000 samples), GenVideo (demamba) (Fake subset) (9,000 samples), DiffusionForensics (dire_wang2023dire) (4,500 samples), WildFake (wildfake_2025) (12,000 samples), and synthetic data generated by Qwen-Image (wu2025qwenimagetechnicalreport) (634 samples). In total, the dataset used for training comprises 106K samples.
Generated Sources
To further enrich the dataset and better capture synthetic images prevalent in real-world scenarios, additional samples were generated using a diverse set of state-of-the-art image generation models. These models span multiple architectures and training paradigms, ensuring broad coverage of visual styles and generative patterns for robust evaluation. Specifically, extra sources includes samples generated by FLUX.1 (labs2025flux1kontextflowmatching), Lora Flux (labs2025flux1kontextflowmatching) and Qwen-Image (wu2025qwenimagetechnicalreport).
3.2 Data Annotation
As illustrated in Table 1, “file_type” indicates the modality of the input—either “image” or “video”, and label represents the ground-truth label, assigned as either “real” or “fake”. The distilled explanations were further categorized along two major dimensions to facilitate structured analysis: Spatial Features, which comprises eight sub-dimensions and captures artifacts and inconsistencies observable within individual frames or static images. Temporal Features, which includes four sub-dimensions and describes anomalies associated with motion, temporal coherence, and cross-frame consistency. Since still images inherently lack temporal attributes, this category is exclusively applicable to video annotations. These categories were informed by established taxonomies of visual artifacts in generative content, as detailed in prior research (deng2024survey).
3.3 Comparison with Existing Datasets
A comparative overview of Ivy-Fake and several existing AIGC detection datasets is provided in Table 2. Notably, Ivy-Fake offers unique advantages by integrating explainable annotations across both image and video modalities, addressing a significant gap in current resources.
Table 2: Comparison on the different datasets used in binary classification and interpretability tasks. Token lengths are computed using the GPT-4o tokenizer from the tiktoken library.
Dataset Avg Token Generator Modality Dataset
Lengths fake real
FakeBench (liu2024mmfakebench) - 6 Image 3K 3K
VANE-Bench (bharadwaj2024vanebench) 101 5 Image* 2K 1K
LOKI (loki_ye2024loki) 226.5 $\sim$16 Image+Video 3K 0
FakeClue (fakeclue_wen2025spot) 119.6 26 Image 68K 36K
Ivy-Fake 451.7 >30 Image+Video 53K 78K
We identified a long-tail distribution in the distilled dataset, with a small portion of samples exhibiting token lengths concentrated at both extremes (xin2026hytrechybridtemporalawareattention). To mitigate the influence of these outliers, we retained only the central 90% of the distribution, filtering out the top and bottom 5% of samples, thereby keeping data with token lengths between 60 and 800. As shown in Figure 2 panel (a), Ivy-Fake exhibits the longest token length distribution, with an average token length of 451.7. Additionally, to conduct a fine-grained analysis of the semantic dimensions embedded in the explanatory language across different interpretability datasets, we employ Qwen3-32B (qwen3) for dimension extraction. To enhance the accuracy of the extraction, we embed supplementary contextual information from each dataset into the model input. For example, in LOKI (loki_ye2024loki), the bounding box coordinates are wrapped with <bbox></bbox> tags, while metadata such as coordinates, titles, and fine-grained descriptions are also included as input to the model. Moreover, we set the Temperature to 0.8 and perform three extraction passes for each sample, retaining only the dimensions consistently identified across all runs. The final aggregated results are visualized in panel (b).
Figure 3: Comparison between Ivy-Fake and FakeVLM (fakeclue_wen2025spot) (NeurIPS 2025).
Ivy-Fake outperforms FakeVLM in terms of interpretability. While FakeVLM only explains superficial factors such as lighting and texture, Ivy-Fake considers multiple semantic dimensions, including hair, skin, and background, leading to a more comprehensive understanding of forged content.
3.4 Quality Control
To mitigate the impact of hallucinations from Gemini 2.5 Pro during question generation, we sampled 8,000 cases for manual review. Each chain-of-thought (CoT) involving Gemini underwent at least two rounds of human verification. A total of ten users, including senior university students and regular participants, contributed to the verification process, which required approximately 1,000 hours in total. Each case was tested by at least two users to ensure robustness across different modalities of synthetic data. After this rigorous filtering, we retained 5,000 cases as the benchmark test set, consisting of 2,500 image samples and 2,500 video samples, with each modality containing 1,250 real and 1,250 fake instances.
3.5 Method
Our preliminary investigations revealed that existing MLLMs exhibit inadequate performance on these tasks. To overcome this limitation, we propose Ivy-xDetector, a multimodal large language model designed explicitly for robust and explainable AIGC detection. The overall training pipeline is illustrated in Figure 4.
Figure 4: Overview of the three-stage training pipeline for Ivy-xDetector.
Stage 1: Cold Start for Instruction-Driven Detection and Explainability.
Due to limited reasoning ability, existing models struggle to detect AIGC content and generate reliable explanations. To address this, we inject Gemini-generated detection and explanatory COT into the training process, thereby improving the model’s fine-grained artifact perception and explanation quality. This stage aims to initialize Ivy-xDetector with instruction following and explainability capabilities. The resulting model not only performs accurate AIGC detection but also generates coherent, human-understandable rationales for its classifications.
Table 3: Performance comparison of models on image and video tasks. “Auto Metrics” include Acc, F1. “GPT Assisted” includes four subjective criteria: Comprehensiveness, Relevance, Detail, and Explanation. Bold indicates the best result, and underline indicates the second best.
Model Image Video Overall
Auto Metrics GPT Assisted Auto Metrics GPT Assisted Auto Metrics GPT Assisted
Acc/F1/ROUGE-L/SIM Com./Rel./Det./Exp. Acc/F1/ROUGE-L/SIM Com./Rel./Det./Exp. Acc/F1/ROUGE-L/SIM Com./Rel./Det./Exp.
Closed-source MLLMs
GPT-4o 0.725/0.723/0.108/0.525 2.34/3.20/2.04/3.26 0.448/0.579/0.072/0.451 1.79/2.35/1.67/2.40 0.587/0.663/0.090/0.488 2.07/2.78/1.85/2.83
Gemini-2.5-Flash 0.747/0.737/0.263/0.733 3.94/4.11/4.04/4.09 0.810/0.811/0.246/0.723 4.00/4.37/4.03/4.36 0.779/0.776/0.254/0.728 3.97/4.24/4.04/4.22
Open-source MLLMs
7B-Parameters Models
InternVL3.5-8B 0.605/0.602/0.194/0.680 2.83/3.49/2.69/3.32 0.574/0.588/0.188/0.664 2.75/3.35/2.68/3.28 0.589/0.596/0.191/0.672 2.79/3.42/2.69/3.30
MiMo-VL-7B 0.662/0.637/0.121/0.593 1.99/2.80/1.85/2.89 0.778/0.783/0.112/0.580 2.04/2.91/1.90/3.19 0.720/0.715/0.116/0.586 2.01/2.86/1.87/3.04
Qwen2.5-VL-7B 0.013/0.026/0.006/0.264 1.02/1.03/1.01/1.52 0.092/0.159/0.015/0.280 1.14/1.23/1.12/2.21 0.053/0.096/0.010/0.272 1.08/1.13/1.07/1.86
LLaVA-OneVision-1.5-8B 0.500/0.333/0.080/0.499 1.51/2.62/1.49/2.38 0.500/0.333/0.068/0.481 1.49/2.26/1.37/2.19 0.500/0.333/0.074/0.490 1.50/2.44/1.43/2.28
MiniCPM-V-4.5 0.666/0.680/0.169/0.637 3.20/3.93/3.06/3.60 0.491/0.505/0.152/0.627 2.83/3.66/2.76/3.36 0.579/0.610/0.161/0.632 3.01/3.80/2.91/3.48
3B-Parameters Models
Qwen2.5-VL-3B 0.641/0.612/0.023/0.391 1.19/1.33/1.18/3.28 0.689/0.686/0.017/0.381 1.42/1.56/1.40/3.68 0.665/0.652/0.020/0.386 1.31/1.45/1.29/3.48
Gemma-3-4b-it 0.408/0.477/0.170/0.576 2.55/3.36/2.46/3.11 0.396/0.482/0.149/0.561 2.30/3.03/2.37/2.95 0.402/0.482/0.159/0.568 2.43/3.19/2.42/3.03
InternVL3.5-2B 0.602/0.573/0.177/0.648 2.62/3.29/2.51/3.13 0.435/0.459/0.159/0.631 2.46/3.16/2.42/2.99 0.518/0.518/0.168/0.640 2.54/3.22/2.47/3.06
InternVL3.5-4B 0.651/0.652/0.190/0.660 3.01/3.68/2.95/3.58 0.614/0.617/0.181/0.653 2.93/3.61/2.83/3.52 0.632/0.635/0.186/0.656 2.97/3.64/2.89/3.55
Ivy-xDetector 0.831/0.831/0.283/0.714 3.54/4.04/3.61/3.85 0.897/0.897/0.300/0.726 3.72/4.12/3.75/4.24 0.864/0.864/0.291/0.720 3.63/4.08/3.68/4.05
Stage 2: Sparse Rewards in Fine-grained Visual forgery Reasoning via Reinforcement Learning
Although the fine-tuned model from Stage 1 demonstrates improved artifact awareness and interpretability, its generalization ability remains limited. To further enhance the model’s capacity for producing consistent and human-understandable explanations, we adopt the Group Relative Policy Optimization (GRPO) algorithm (shao2024deepseekmathpushinglimitsmathematical). GRPO enables reinforcement learning (zhoubao-earlvr; zhao2026know) using only binary classification samples, achieving efficient and data-light optimization.
We construct binary real/fake pairs from multiple datasets, each containing both authentic and synthetic images and videos. Ivy-xDetector takes the visual input and extracts the text enclosed within the <conclusion></conclusion> tags, where the predicted label is either real or fake. The quantitative results of different training stages are presented in Table 7.
3.6 GRPO
Following DeepSeek-R1 (deepseekr1_guo2025deepseek) and TextShield-R1 qu2026textshield, we adapt the Group Relative Policy Optimization (GRPO) (shao2024deepseekmathpushinglimitsmathematical), an online RL algorithm designed to maximize the advantage of generated completions while constraining policy divergence from a reference model. We formalize our training process of Ivy-xDetector using GRPO below. Let $p$ denote a sampled prompt, and let $o_{1},o_{2},...,o_{n}$} be a group of completions generated by the current policy $\pi_{\theta}$. For each completion $G_{i}$, a reward $r_{i}$ is computed using a rule-based reward function. The group advantage for each completion is then calculated as:
$\hat{A}_{i,t}=\frac{r_{i}-mean(r)}{std(r)}$ (1)
$\mathcal{L}_{\mathrm{GRPO}}(\theta)=-\frac{1}{K}\sum_{k=1}^{K}A^{(k)}\,\ell^{(k)}(\theta)\;+\;\beta\,\mathrm{KL}\!\left(\pi_{\theta}(\cdot|x)\,\|\,\pi_{\mathrm{ref}}(\cdot|x)\right).$ (2)
where $\beta$ is a coefficient that balances advantage maximization and KL regularization, and the clipping operator $clip(\dots,1-\epsilon,1+\epsilon)$ constrains the update magnitude. To regularize policy updates, we estimate the token-level Kullback-Leibler (KL) divergence between the current policy $\pi_{\theta}$ and a reference policy $\pi_{ref}$.
$\mathbb{D}_{\text{KL}}\big[\pi_{\theta}\big\|\pi_{\text{ref}}\big]=\frac{\pi_{\text{ref}}(o_{i,t}\mid p,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid p,o_{i,<t})}-\log\frac{\pi_{\text{ref}}(o_{i,t}\mid p,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid p,o_{i,<t})}-1.$ (3)
3.7 Reward Model
For effective RL, we employ a rule-based reward that consists of accuracy and format rewards. We introduce a solid accuracy reward for AIGC Detection, which utilizes distinct functions to evaluate binary classification task. This allows for a more appropriate assessment based on the expected answer type.
  • Accuracy Reward: The accuracy reward assigns a score of 1 if the label in <conclusion> exactly matches the ground-truth classification real and fake and 0 otherwise.
  • Format Reward: The format reward assigns a score of 1 if the output strictly follows the structural requirements by enclosing the reasoning within <think></think> tags and the final decision within <conclusion></conclusion> tags, and 0 otherwise.
4 Experiments
4.1 Experimental Details
Baselines
We primarily evaluate three closed-source models and three open-source models on our Ivy-Fake. The closed-source models are GPT-4o (gpt4_achiam2023gpt) and Gemini2.5 Flash (genimi_team2023gemini). For the open-source models, we select models of comparable size: InternVL3.5 (chen2024internvl), Qwen2.5-VL (bai2025qwen2), MiMo-VL (coreteam2025mimovltechnicalreport), LLaVA-OneVision-1.5 (LLaVA-OneVision-1.5; lillava) and MiniCPM-V-4.5 (yao2024minicpm; yu2025minicpmv45cookingefficient). Ivy-xDetector denotes our fully RL-trained model after the GRPO optimization stage.
Evaluation
We report standard accuracy (Acc) and macro-averaged F1 score (F1) to assess the model’s ability to distinguish real from fake instances. For the reasoning task, we measure the similarity between the model’s reasoning process and the reference annotations using the ROUGE-L score (lin2004rouge) and BertScore (bert-score; he2021deberta), which captures the longest common subsequence between predicted and reference texts, reflecting token-level overlap. Since ROUGE-L may fail to fully capture the fidelity of reasoning steps, we adopt an LLM-as-a-judge evaluation paradigm (zheng2023judging), following the FakeBench protocol (liu2024mmfakebench), which assesses model responses along four dimensions: (1) Completeness: It reflects the extent to which the response fully addresses all aspects of the user’s question. More complete responses should incorporate information aligning well with the “golden clues” or reference answers. Incomplete or partially answered responses will receive lower scores. (2) Relevance: Measure how closely the content relates to the original annotation; (3) Level of Detail: Assess whether the response includes enough examples or elaborations; (4) Explanation: Verify the accuracy and consistency of explanations for any causes mentioned. Each response is scored using GPT-4o mini (gpt4_achiam2023gpt), which instructs the model to act as an impartial judge and assign a score from 1 to 5.
Figure 5: Ivy-xDetector performance on video detection.
Implementation Details
Our model was trained using the AdamW (adamw_Loshchilov2017FixingWD) optimizer with a cosine learning rate schedule. During the SFT stage, the model was trained for one epoch with a batch size of 1, which ensures stable convergence on large-scale multimodal data. In the subsequent RL stage, we adopted the GRPO algorithm, training for one epoch with the same batch size and a learning rate of $1\times 10^{-5}$. Within GRPO, the group size $n$ was set to 8, the warm-up ratio to 0.01, the temperature to 0.9. Training the SFT and RL stages required approximately 10 and 50 hours, respectively, on a system equipped with 32 A100 GPUs.
For video-based inputs, we configured the preprocessing pipeline to sample frames at a rate of 1 frame per second, with a maximum of 6 frames per clip to ensure temporal coherence while maintaining computational efficiency. Across the training pipeline, the overall input resolution was constrained by a global upper bound of 6,422,528 pixels (MAX_PIXELS), preventing overflow during multi-GPU parallel training and ensuring consistent input scaling across image and video modalities.
4.2 Main Results
Table 4: Comparison on the Chameleon (aide_yan2025a). Accuracy (%) of different detectors (rows) in detecting real and fake images. For each training dataset, the first row is overall accuracy, the second row is “fake/real” class accuracy.
CNNSpot FreDect Fusing GramNet LNP UnivFD DIRE PatchCraft NPR AIDE Ivy-xDetector
60.89 57.22 57.09 60.95 58.52 60.42 59.71 56.32 58.13 65.77 73.17
9.86/99.25 0.89/99.55 0.02/99.98 4.76/99.66 7.72/96.70 85.52/41.56 11.86/95.67 3.07/96.35 2.43/100.00 26.80/95.06 67.78/77.22
Table 5: Comparison on the Genimage (genimage_2023). Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). The best result and the second-best result are marked in bold and underline, respectively. Results of other methods are reported from (aide_yan2025a).
Method Midjourney SD v1.4 SD v1.5 ADM GLIDE Wukong VQDM BigGAN Mean
CNNSpot (cnnspot_wang2020cnn) 52.80 96.30 95.90 50.10 39.80 78.60 53.40 46.80 64.21
F3Net (qian2020thinking) 50.10 99.90 99.90 49.90 50.00 99.90 49.90 49.90 68.69
DIRE (dire_wang2023dire) 60.20 99.90 99.80 50.90 55.00 99.20 50.10 50.20 70.66
GenDet (zhu2023gendet) 89.60 96.10 96.10 58.00 78.40 92.80 66.50 75.00 81.56
PatchCraft (aigcdetectionbenchmark_zhong2023patchcraft) 79.00 89.50 89.30 77.30 78.40 89.30 83.70 72.40 82.30
FRIDA (bonechi2025thisfakedetectionsource) 85.80 85.00 85.50 84.00 87.30 83.90 86.50 84.80 85.30
AIDE (aide_yan2025a) 79.38 99.74 99.76 78.54 91.82 98.65 80.26 66.89 86.88
DRCT (drct) 91.50 95.00 94.40 79.40 89.20 94.70 90.00 81.70 89.50
Effort (effort) 82.40 99.80 99.80 78.70 93.30 97.40 91.70 77.60 91.10
ThinkFake-7B (huang2025thinkfakereasoningmultimodallarge) 92.50 93.10 95.30 73.10 87.40 93.60 66.20 70.80 84.00
Ivy-xDetector 96.41 97.07 97.14 95.59 95.95 96.92 96.03 95.22 96.32
Table 6: Comparisons to the GenVideo (demamba). F1 score (F1), recall score (R) and average precision (AP) on the many-to-many generalization task. “Demamba-XCLIP-FT” is abbreviated as “Demamba”. Results of other methods are reported from (demamba).
Model Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Avg.
Studio Valley Scope Scrape
F3Net (Image) R 0.8393 0.9971 0.9862 0.7757 0.5700 0.3657 0.9952 0.9971 0.8943 0.7678 0.8188
F1 0.5000 0.9406 0.9628 0.8169 0.6988 0.4904 0.9332 0.9688 0.8873 0.8251 0.8024
NPR (Image) R 0.9107 0.9957 0.9949 0.2429 0.8964 0.5771 0.9712 0.9986 0.9429 0.8780 0.8408
F1 0.2786 0.8441 0.9131 0.3028 0.8627 0.5944 0.8170 0.9164 0.8184 0.8163 0.7164
STIL (Video) R 0.7857 0.9814 0.9804 0.7600 0.6179 0.5329 0.9936 0.9736 0.9457 0.6501 0.8222
F1 0.3805 0.9068 0.9458 0.7824 0.7232 0.6217 0.9039 0.9433 0.8884 0.7267 0.7823
DeMamba (Video) R 0.9812 1.0000 0.9986 0.6543 0.9486 0.9886 1.0000 1.0000 0.9286 0.8909 0.9302
F1 0.6407 0.9602 0.9790 0.7539 0.9537 0.9551 0.9557 0.9797 0.9240 0.9120 0.9020
Ivy-xDetector R 0.7857 0.9371 0.9507 0.9443 0.9550 0.9643 0.9968 0.9857 0.8943 0.9461 0.9528
F1 0.8800 0.9676 0.9747 0.9713 0.9770 0.9818 0.9984 0.9928 0.9442 0.9723 0.9526
We perform extensive experiments to assess both detection and explanation capabilities. In particular, the proposed method is evaluated on the classification (real/fake) tasks and reasoning tasks for both image and video content using the proposed unified framework. For the classification task, we test our model on both image and video content to detect the synthesic content.
From Table 3, our method demonstrates robust detection and explanation capabilities across both image and video tasks. Overall, closed-source models maintain a clear advantage in classification accuracy and subjective explanation quality. Among them, Gemini-2.5-flash achieves the best performance on both automatic metrics and GPT-assisted evaluation (e.g., reaching Acc/F1 of 0.812/0.812 on video detection and an average explanation score above 4.0), reflecting strong overall detection and reasoning ability. In contrast, GPT-4o maintains relatively high classification accuracy.
For open-source models, most models in the 3B–7B parameter range exhibit relatively low scores on explanation dimensions (<3.5<3.5), indicating limitations in generating reasoning chains and covering multi-dimensional details. Notably, Qwen2.5-VL-3B shows relatively stable performance in explanation relevance and detail (Det./Exp. >3.2>3.2), suggesting that lightweight models still hold potential under specific designs.
Notably, Gemini-2.5-Flash achieves the highest scores across all explanation-related dimensions, with an average of 4.12, surpassing all other models. This advantage can be attributed to the fact that part of our training data was distilled from Gemini-2.5-Pro, a powerful multimodal model capable of handling complex reasoning and explanation tasks. In contrast, our Ivy-xDetector ranks second with an average explanation score of 3.85, demonstrating interpretability comparable to Gemini-2.5-Flash while achieving substantially higher detection accuracy (Acc/F1 = 0.864/0.864) than any other open-source model. Compared with Gemini-2.5-Flash (Acc/F1 = 0.780/0.776), Ivy-xDetector improves accuracy by over 10 percentage points, confirming its robustness in both synthetic image detection and interpretive reasoning tasks.
4.3 Generalization Evaluation
Evaluation on Classification Benchmarks.
We conduct evaluations across multiple public leaderboards, including the image-based GenImage 5 and the video-based GenVideo benchmark 6. Performance results of other competing detectors are taken from AIDE (aide_yan2025a) and Demamba (demamba). Our method achieves superior performance on all these leaderboards, demonstrating superior detection accuracy and robust cross-modal consistency.
Evaluation on Unseen Benchmarks.
Beyond the datasets aligned with our training sets, we further evaluate the generalization ability of Ivy-xDetector on the Chameleon benchmarkaide_yan2025a, which lies outside the training data distribution, as shown in Table 4. The results confirm that our approach preserves strong generalization capability across unseen generative models and diverse data domains.
Compared to the original leaderboard SOTA method AIDE, our model achieves an overall accuracy of 73.17%. Notably, while the accuracy for the real class decreases from 95.06% to 77.22%, the accuracy for the fake class substantially increases from 26.80% to 67.78%. This suggests that AIDE tends to overfit the real category, whereas our model achieves a more balanced and robust performance across both real and synthetic content.
Ablation Study.
Table 7: Ablation study on different SFT and RL training settings. Base Model denotes the base model Qwen2.5-VL-3B. Evaluation metric: ACC / F1.
Model Name Training Stage Chameleon (Image) GenVideo (Video) Ivy-Fake Test
SFT RL
Base Model 59.05 / 53.30 67.70 / 80.74 65.10 / 61.80
- 40.31 / 37.99 91.61 / 91.55 73.91 / 73.83
- 72.72 / 70.79 78.19 / 77.62 77.00 / 76.66
- 73.17 / 73.14 95.29 / 95.29 86.40 / 86.30
To validate the effectiveness of our training framework, we design four ablation variants, as shown in Table 7. As shown, applying SFT alone results in only a slight performance drop on the out-of-domain benchmark Chameleon, while achieving consistent improvements across all other datasets. This suggests that the SFT stage primarily enhances the model’s fine-grained sensitivity to synthetic artifacts. In contrast, applying RL alone yields modest but consistent gains in classification accuracy across benchmarks. The combination of SFT and RL achieves the best balance between accuracy and interpretability.
5 Conclusion
We introduced Ivy-Fake, the first unified, large-scale dataset for explainable AIGC detection across both images and videos, featuring over 106K richly annotated training samples and 5,000 evaluation examples with natural-language reasoning, and proposed Ivy Explainable Detector (Ivy-xDetector), a MLLM that jointly detects and explains synthetic content. Our model sets superior benchmarks in AIGC detection and explainability, and our publicly released resources provide a robust foundation for transparent, trustworthy multimodal analysis. We believe that such fine-grained explainability data Ivy-Fake can catalyze new research directions for the AIGC detection community, enabling deeper understanding of synthetic content and more principled model development.
Limitations: Future work will concentrate on tailoring the benchmark toward domain-specific scenarios, including explainability-focused datasets for DeepFake and document manipulation analysis. Nevertheless, our experiments suggest that the proposed approach exhibits strong scalability and generalization capabilities, indicating the potential for even better performance when applied to larger models such as 32B or 72B.
Broader impacts: The interpretability evaluation framework we propose is currently applied only to the field of Fake Image Detection (du2025forensichub), which is strictly a binary classification task. In future work, our multi-dimensional evaluation approach can be extended to assess the interpretability of other tasks, such as chain-of-thought evaluation for Table Question Answering (jiang2025tabdsr), question answering in medical scenarios (zhao2026spinebench; xin2024artificial; xin2026dualcpt), GUI Agents (yuanse), Document Generation ying2024dig; ying2024dig; ying2024fine and video grounding interpretability (lin2025audio).